# OptiMind: Teaching LLMs to Think Like Optimization Experts

### v1

@@ 5.1 Experimental Setup
Datasets. Our evaluation is performed on our rigorously cleaned versions of three challenging and widely-used benchmarks: IndustryOR (Tang et al., 2024), Mamo-Complex (Huang et al., 2024), and OptMATH (Lu et al., 2025), each with around 100-160 problems (see Appendix A.9 for details). We selected these datasets as they are commonly used in previous works and represent some of the most complex formulation tasks in the literature, providing a strong signal for model capabilities. While we also considered other benchmarks such as OptiBench (Yang et al., 2025) (605 problems) and ComplexOR (Xiao et al., 2024) (18 problems), our initial experiments revealed that performance on these datasets has either saturated or does not effectively differentiate between models of varying scales; moreover, their sizes are either too small (yielding a noisy evaluation signal) or too large to feasibly clean with the same rigor. A more detailed discussion of our dataset selection and the comprehensive cleaning process is provided in Appendix A.8.
Metrics. Our primary metric is solution accuracy, which we report at the first generation turn (Turn 1) and after five turns of iterative self-correction (Turn 5). Our prompts require an executable Python script using GurobiPy to formulate and solve the problem. Following SIRL (Chen et al., 2025), a solution is correct if its objective matches the ground truth within both relative and absolute error of $10^{-6}$. To extract this value, we insert a print statement emitting the optimal objective with a unique tag and parse it from the execution log, as in SIRL. We also assess the effect of self-consistency by comparing results without majority voting ($K=1$) against results with majority voting over $K=8$ samples, grouping answers within a relative and absolute tolerance of $10^{-6}$ to account floating-point variations. To ensure statistical robustness, all reported results are averaged across 10 independent experiments using different random seeds.
@@ 5.2 Main Results
Figure 5: Average accuracy for different models (without majority voting). Our model outperforms the open-source models, and performs comparably to the larger GPT-5 model.
Table 1: Accuracies under different base models and hint settings on expert-cleaned benchmarks.
@@
gpt-5 1 no $78.57\pm 3.09$ $81.12\pm 1.79$ $87.86\pm 1.55$ $89.36\pm 1.92$ $90.15\pm 0.69$ $92.18\pm 0.78$
yes $83.50\pm 2.84$ $85.20\pm 1.25$ $92.57\pm 1.19$ $93.42\pm 0.62$ $89.21\pm 1.78$ $90.31\pm 1.04$
o4-mini 1 no $76.19\pm 1.05$ $82.14\pm 2.11$ $75.95\pm 2.38$ $87.78\pm 1.15$ $80.31\pm 2.02$ $87.18\pm 0.89$
yes $82.14\pm 2.47$ $5.88\pm 1.19 88$$8.09\pm 2.02$ $91.67\pm 1.52$ $84.21\pm 2.17$ $90.46\pm 1.16$
Qwen3-32B 1 no $64.28\pm 3.00$ $9.39\pm 3.00 66$$6.47\pm 1.64$ $77.14\pm 3.18$ $79.01\pm 2.41$ $79.49\pm 1.84$
yes $69.13\pm 2.00$ $74.23\pm 3.85$ $78.00\pm 1.73$ $84.76\pm 2.12$ $76.95\pm 1.35$ $80.85\pm 1.35$
SIRL-Gurobi32B 1 no $1.36\pm 1.39 55$$5.27\pm 1.63$ $64.00\pm 1.41$ $72.67\pm 1.37$ $70.93\pm 1.94$ $75.46\pm 1.62$
yes $49.38\pm 3.02$ $53.06\pm 3.61$ $75.61\pm 2.05$ $78.57\pm 0.95$ $74.06\pm 1.39$ $76.71\pm 2.30$
gpt-oss-20b 1 no $67.14\pm 3.46$ $73.67\pm 2.88$ $70.00\pm 1.93$ $86.57\pm 0.91$ $76.32\pm 2.63$ $86.32\pm 2.12$
yes $\bf{72.45\pm 1.66}$ $\bf{76.83\pm 1.52}$ $83.81\pm 3.05$ $89.90\pm 1.73$ $79.84\pm 2.04$ $86.56\pm 1.75$
Our SFT 1 no $68.36\pm 2.92$ $74.92\pm 2.42$ $74.42\pm 1.15$ $86.05\pm 1.55$ $80.35\pm 3.10$ $\bf{87.94\pm 1.41}$
yes $70.26\pm 1.90$ $74.20\pm 2.61$ $\bf{85.85\pm 0.76}$ $\bf{90.41\pm 0.58}$ $\bf{81.13\pm 1.59}$ $87.39\pm 1.05$
gpt-oss-20b 8 no $76.22\pm 3.15$ $78.37\pm 2.49$ $6.09\pm 3.55 88$$8.76\pm 1.72$ $86.95\pm 3.08$ $89.53\pm 2.27$
yes $\bf{80.10\pm 2.69}$ $79.69\pm 2.98$ $90.04\pm 2.76$ $90.72\pm 1.08$ $\bf{88.36\pm 2.06}$ $\bf{89.14\pm 1.40}$
Our SFT 8 no $75.85\pm 1.56$ $79.25\pm 0.58$ $7.61\pm 0.89 88$$8.19\pm 0.78$ $7.76\pm 5.31 88$$8.28\pm 4.00$
yes $78.91\pm 2.57$ $\bf{80.91\pm 1.77}$ $\bf{92.14\pm 0.61}$ $\bf{92.02\pm 0.71}$ $7.76\pm 1.19 88$$8.28\pm 2.06$
We now assess the impact of the components of OptiMind’s pipeline on solution accuracy. Accuracy improvements are reported as absolute percentage points.
Results with OptiMind’s training data cleaning. First, SFT training fed with the cleaned training data processed through OptiMind has demonstrable gains. Focusing on the gpt-oss-20b model and isolating training effects under the single-turn without hints, our SFT model outperforms the base model by $1.2\%$ on IndustryOR, $4.4\%$ on Mamo-Complex, and $4.0\%$ on OptMATH. Improvement gains remain under 5-turn self-correction setting (see Table 2 for details). Additional ablation study can be found in §5.3.
Table 2: GPT-OSS-20B vs. our SFT model under no-hint prompting, comparing single-turn ($K=1$) and 5-turn self-correction ($K=1$). We report SFT improvements over the base in percentage points.
Benchmark Base (1 turn) SFT (1 turn) Gain (%) Base (5 turns) SFT (5 turns) Gain
IndustryOR 67.14 68.36 +1.22 73.67 74.92 +1.25
Mamo-Complex 70.00 74.42 +4.42 86.57 86.05 -0.52
OptMATH 76.32 80.35 +4.03 86.32 87.94 +1.62
Results with OptiMind’s error analysis. Additionally, using class-specific error analysis and the associated hints at inference consistently lift single-turn accuracy across models and datasets. Typical gains from “no hint” to “with hints” across the baselines and our SFT model (single-turn, no majority voting) range from are $3\%$ to $6\%$ on IndustryOR, $11\%$ to $14\%$ on Mamo-Complex, and $1\%$ to $4\%$ on OptMATH, with rare small dips around $1\%$ to $2\%$ on OptMATH for a few models. Interestingly, even GPT-5 sees a significant improvement in accuracy when using the hints (up to $4.93\%$ on IndustryOR), suggesting that hints can enhance the performance of even very strong models, encoding domain-specific information that seems complementary to general model’s capabilities. See Fig. 6 for the detailed results. As a deeper dive, we examine Mamo-Complex, which is dominated by five problem classes. Fig. 7 (left) shows a per-class breakdown of gpt-oss-20b model without and with hints at inference: accuracy increases are broad and most pronounced on types prone to error related to sign conventions (e.g., within flow-conservation-type constraints) and in complex structural constraints (e.g., subtour elimination constraints in TSP). While IndustryOR and OptMATH contain more heterogeneous types, we observe the same consistent pattern of improvement; see Appendix A.7 for details.
Overall comparison. Our results are summarized in Table 1. As can be seen, the model obtained with the complete OptiMind pipeline, that is, gpt-oss-20b model with SFT, error-analysis hints, multi-turn inference and majority voting, attains accuracy that is close to much larger state-of-the-art models and clearly outperforms open-source models of similar size. For example, when comparing against GPT-5 (single turn, no hints), our model is within $1.9\%$ on OptMATH and is $2.3\%$ and +4.2+4.2 on IndustryOR and Mamo-Complex respectively. In addition, we see sizable improvements over the base gpt-oss-20b model, even when it uses majority voting: $1.3\%$, $4.7\%$, and $5.9\%$ on OptMATH, IndustryOR, and Mamo-complex, respectively. Other OSS models fall significantly behind ours; for example, comparing with SIRL-Gurobi-32B, we obtain improvements of $17\%$$28\%$ across the different benchmarks.
@@
Figure 7: Left: Accuracy by problem type on Mamo-Complex for gpt-oss-20b (single-turn, no majority voting). Infrequent problem types are omitted. Right: Weighted mean accuracy over five self-debugging turns on the three benchmarks (each weighted by its sample count).
@@ 5.3 Additional Ablation Studies
Ablation on training data. To assess whether our data cleaning improves training, we train three SFT gpt-oss-20b models using progressively higher data quality. (1) Original data: the original OptMATH set plus COPTPY-to-Python/GurobiPy translations from o4-mini, retaining instances whose final solutions match the provided answers; (2) Answer-aligned data: we keep the original problem and final answer, generate 64 solutions with gpt-oss-20b, and keep the sample that matches the original answer even if the answer is imperfect. (3) Cleaned data (ours): we apply our full cleaning pipeline described by §4.3. Table 3 compares single-turn, no-hint accuracy across the three checkpoints. Performance on the original data is poor and, likely due to distribution mismatch in reasoning and lack of reliable thought trajectories. Training on answer-aligned data helps but remains below our cleaned set, suggesting that matching potentially incorrect answers reinforces systematic mistakes and brittle patterns.
Table 3: Single-turn accuracies of SFT variants on different data quality. Numbers are reported by percentage points.
SFT Dataset Hints IndustryOR Mamo-Complex OptMATH
Original Data no $44.48\pm 4.81$ $3.90\pm 2.41 55$$5.31\pm 2.27$
yes $43.36\pm 2.70$ $61.23\pm 2.19$ $80.56\pm 3.99$
Answer-Aligned Data no $66.90\pm 2.62$ $67.75\pm 2.08$ $80.35\pm 1.30$
yes $68.58\pm 2.15$ $84.55\pm 1.90$ $80.24\pm 1.67$
Cleaned Data (ours) no $68.36\pm 2.92$ $74.42\pm 1.15$ $80.35\pm 3.10$
yes $70.26\pm 1.90$ $85.85\pm 0.76$ $81.13\pm 1.59$
Ablation on multi-turn inference. Fig. 7 (right) plots the weighted average accuracy over the number of turns for the base gpt-oss-20b and our SFT model, under the setting of with and without hints. As expected, the accuracy increases monotonically with more turns for both models; however there are diminishing returns that may be taken into account under compute/time vs. accuracy tradeoffs. Overall, the combination SFT + hints is the most effective across benchmarks, especially for small number of turns.
Ablation on Majority Voting (MV). We evaluate the utility of MV when multi-turn correction with tool feedback is available. We perform this ablation on the base gpt-oss-20b model, with and without multi-turn inference. Table 4 shows the effect of MV for the different benchmarks. The main takeaways are: (i) MV and multi-turn are complementary, as both yield sizable gains; (ii) the benefits of MV considerably diminish when multiple turns are applied.
Table 4: Accuracy of gpt-oss-20b across majority voting $K$ and turn counts; hints provided in the prompt. Numbers are reported in percentage points.
@@
K 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns
1 72.45{72.45} 76.83{76.83} $83.81$ $89.90$ $79.84$ $86.56$
8 80.10{80.10} $79.69$ $90.04$ $90.72$ 88.36{88.36} 89.14{89.14}
6 Conclusion
We present OptiMind – a framework for formulating mixed-integer linear optimization problems with LLMs, combining structured data cleaning with error-aware prompting, self-consistency, and tool-driven multi-turn correction. Supervised fine-tuning on cleaned data aligns the model with optimization primitives, while targeted hints improve reliability at inference time. The resulting pipeline attains strong performance across multiple benchmarks. While future LLMs may embed more expert knowledge, we believe the principles and techniques of our framework will remain essential. They can be applied to domains such as supply chain management or adapted to enterprise-specific scenarios to drive real-world impact.
Appendix A Appendix
A.1 Sample Problem Description and Output Optimization Model

### v2

@@ 5.1 Experimental Setup
Datasets. Our evaluation is performed on our rigorously cleaned versions of three challenging and widely-used benchmarks: IndustryOR (Tang et al., 2024), Mamo-Complex (Huang et al., 2024), and OptMATH (Lu et al., 2025), each with around 100-160 problems (see Appendix A.10 for details). We selected these datasets as they are commonly used in previous works and represent some of the most complex formulation tasks in the literature, providing a strong signal for model capabilities. While we also considered other benchmarks such as OptiBench (Yang et al., 2025b) (605 problems) and ComplexOR (Xiao et al., 2024) (18 problems), our initial experiments revealed that performance on these datasets has either saturated or does not effectively differentiate between models of varying scales; moreover, their sizes are either too small (yielding a noisy evaluation signal) or too large to feasibly clean with the same rigor. A more detailed discussion of our dataset selection and the comprehensive cleaning process is provided in Appendix A.9. For completeness, we also report results on cleaned versions of IndustryOR released by SIRL (Chen et al., 2025), LogiOR (Yang et al., 2025a), and Xiao et al. (2025a), as well as on cleaned versions of Mamo-Complex from SIRL and Xiao et al. (2025a).
Metrics. Our primary metric is solution accuracy, which we report at the first generation turn (Turn 1) and after five turns of iterative self-correction (Turn 5). Our prompts require an executable Python script using GurobiPy to formulate and solve the problem. Following OptMATH (Lu et al., 2025), we mark a solution as correct if the normalized error $|\hat{z}-z^{\star}|/(|z^{\star}|+1)$ is below $10^{-6}$, where $\hat{z}$ is the model’s objective value and $z^{\star}$ is the ground-truth objective. To extract this value, we insert a print statement emitting the optimal objective with a unique tag and parse it from the execution log, as in SIRL. We also assess the effect of self-consistency by comparing results without majority voting ($K=1$) against results with majority voting over $K=8$ samples, grouping answers within a relative and absolute tolerance of $10^{-6}$ to account floating-point variations. To ensure statistical robustness, all reported results are averaged across 10 independent experiments using different random seeds.
@@ 5.2 Main Results
Figure 5: Average accuracy for different models without majority voting. Models colored in blue are open-sourced models under 32B parameters, and models in purple are closed-sourced frontier models. OptiMind outperforms open-source baselines and performs comparably to O4-Mini in the single-turn setting and to GPT-5 with multi-turn self-corrections.
We now assess the impact of OptiMind’s pipeline on solution accuracy.
Overall comparison. Our results are visualized in Figure 5, and the full numerical values are reported in Table 2. As can be seen, OptiMind achieves comparable performance to the frontier close-sourced models with much larger size while clearly outperforming other open-source models of similar size. In the single-turn, no-hint regime, our SFT model consistently improves $4\%$-$23\%$ over the gpt-oss-20b base model, $7\%$-$23\%$ over Qwen3-32B, and $13\%$-$26\%$ over the domain-specific state-of-the-art model SIRL-Gurobi32B across all three benchmarks. Combining SFT with error-analysis hints, our OptiMind framework performs comparably to, and often improves upon o4-mini under the single-turn setting: it’s within $1.9\%$ on IndustryOR, and improves by $13.9\%$ on Mamo-Complex, and $3.8\%$ on OptMATH. With five turn, OptiMind further substantially improve upon o4-mini and perform comparably with GPT-5: it is within about $7\%$ on IndustryOR and $1.8\%$ OptMATH, and exceeds it by $4.2\%$ on Mamo-Complex. We also observe similar trends on the cleaned versions of IndustryOR and Mamo-Complex released by prior work (Chen et al. (2025), Yang et al. (2025a), and Xiao et al. (2025a)); detailed numbers for these additional test sets are reported in Table 3 in Appendix A.1.
Results with OptiMind’s SFT training. Our SFT training fed with the cleaned training data processed through OptiMind has demonstrable gains. Under the single-turn without hints, our SFT model outperforms the base model by $2.7\%$ on IndustryOR, $20.7\%$ on Mamo-Complex, and $9.9\%$ on OptMATH. These gains remain when we apply error-analysis hints: with $2.2\%$, $13.2\%$, and $5.4\%$ on IndustryOR, Mamo-Complex, and OptMATH, respectively, as shown in Figure 6. Additional ablation studies can be found in §5.3.
Results with OptiMind’s error analysis. Additionally, using class-specific error analysis and the associated hints at inference consistently lift single-turn accuracy across models and datasets. Typical gains from “no hint” to “with hints” across the baselines and our SFT model (single-turn, no majority voting) range from $1\%$ to $6\%$ on IndustryOR, $2\%$ to $10\%$ on Mamo-Complex, and $1\%$ to $4\%$ on OptMATH, with rare small dips around $2\%$ on OptMATH for a few models. Interestingly, even GPT-5 sees a significant improvement in accuracy when using the hints (up to $4.71\%$ on MAMO-Complex), suggesting that hints can enhance the performance of even very strong models, encoding domain-specific information that seems complementary to general model’s capabilities. See Fig. 7 for the detailed results.
As a deeper dive, we examine Mamo-Complex, which is dominated by five problem classes. Fig. 8 shows a per-class breakdown of gpt-oss-20b model without and with hints, and and our SFT model with hints at inference: accuracy increases are broad and most pronounced on types prone to error related to sign conventions (e.g., within flow-conservation-type constraints) and in complex structural constraints (e.g., subtour elimination constraints in TSP), and the SFT+hint setting further improves over the base+hint model across these classes. While IndustryOR and OptMATH contain more heterogeneous types, we observe the same consistent pattern of improvement; see Appendix A.8 for details.
Effect of test-time scaling
We evaluate the effect of two test-time scaling techniques—multi-turn self-correction and majority voting (MV)—when additional computation is available. Figure 9 show that both techniques provide consistent accuracy improvements over single-turn inference for the gpt-oss-20b base model and our SFT model, with and without hints. MV and multi-turn are complementary, but the marginal benefit of MV diminishes once multiple turns are used, and gains from additional turns also exhibit diminishing returns as compute increases. Overall, the combination of SFT and hints with a small number of self-correction turns delivers the best accuracy-compute tradeoff across benchmarks.
Figure 6: GPT-OSS-20B vs. our SFT model under no-hint prompting, comparing the prompts without hint and with hint under single-turn, no majority-voting setting. We report SFT improvements over the base in percentage points.
Figure 8: Accuracy by problem type on Mamo-Complex for gpt-oss-20b (single-turn, no majority voting) and our SFT model. Infrequent problem types are omitted.
@@
gpt-oss-20b 1 no $66.23\pm 3.77$ $71.89\pm 3.05$ $65.03\pm 3.91$ $86.80\pm 1.33$ $74.87\pm 3.40$ $86.45\pm 2.13$
yes $68.28\pm 2.61$ $75.05\pm 1.84$ $75.85\pm 2,74$ $89.47\pm 1.27$ $76.95\pm 3.02$ $87.57\pm 1.24$
gpt-oss-20b 8 no $5.85\pm 2.04 77$$7.47\pm 1.26$ $89.71\pm 1.12$ $89.66\pm 0.89$ $90.31\pm 1.17$ $90.39\pm 1.64$
yes $8.48\pm 2.07 77$$7.87\pm 2.25$ $92.00\pm 0.58$ $91.71\pm 0.78$ $91.14\pm 1.10$ $89.32\pm 1.03$
Our SFT 1 no $68.97\pm 2.15$ $72.43\pm 3.43$ $85.77\pm 2.19$ $90.71\pm 1.24$ $\bf{84.82\pm 2.95}$ $\bf{89.62\pm 2.80}$
yes $\bf{70.50\pm 3.29}$ $\bf{75.25\pm 2.57}$ $\bf{89.09\pm 1.27}$ $\bf{92.04\pm 1.05}$ $2.34\pm 1.81 88$$8.35\pm 1.58$
Our SFT 8 no $78.48\pm 1.43$ $79.09\pm 1.58$ $\bf{92.95\pm 0.58}$ $\bf{93.09\pm 0.66}$ $\bf{93.55\pm 0.69}$ $\bf{93.94\pm 1.00}$
yes $\bf{79.39\pm 1.27}$ $\bf{79.89\pm 1.11}$ $92.00\pm 0.58$ $91.71\pm 0.78$ $88.99\pm 0.95$ $89.74\pm 1.06$
@@
Figure 9: Left: Accuracies of the gpt-oss-20b base model and our SFT model across majority voting $K$ and turn counts on the three benchmarks. Right: Weighted mean accuracy over five self-correcting turns on the three benchmarks (each weighted by its sample count). Results for each benchmark can be found in Appendix A.1.
@@ 5.3 Additional Ablation Studies
Ablation on training data. To illustrate the effectiveness of our cleaned training data, we evaluate seven alternative data configurations: (1) Original data: the original OptMATH set plus COPTPY-to-Python/GurobiPy translations from o4-mini, keeping only instances whose final solutions match the provided answers; (2) No OR-Instruct fixing: training on the original OR-Instruct questions without fixing missing-data issues; (3) No OptMATH back-translation: training on the original OptMATH questions without back-translation; (4) Single-shot solution: generating solutions without majority voting; (5) No-hint: removing error-analysis hints when re-generating solutions; (6) Answer-aligned data: keeping the original problem and final answer, generating 64 solutions with gpt-oss-120b, and selecting the sample that matches the original answer even if that answer is imperfect; and (7) 20b data: replacing gpt-oss-120b with gpt-oss-20b as the response generator for training trajectories.
Figure 10 compares these configurations under the single-turn, no-hint setting and reports the weighted average accuracy across the three benchmarks (weighted by sample count). All ablations underperform our full data recipe, which combines missing-data fixes, manual resolution of key ambiguities, OptMATH re-generation with back-translation, and gpt-oss-120b-generated trajectories. In particular, using gpt-oss-20b to generate data)is competitive but consistently weaker than the gpt-oss-120b-based data, showing that our full data recipe is effective.
Figure 10: Ablation on training-data configurations. We compare seven variants of our training data recipe under the single-turn, no-hint setting and report weighted average accuracy across IndustryOR, Mamo-Complex, and OptMATH.
Ablation on classification. We also test the classification component on the Mamo-Complex benchmark, comparing four configurations: (1) classification performed by gpt-oss (our default); (2) (approximate) oracle classification by human expert manually labeled classes; (3) random assignment of classes; and (4) applying all class-specific hints together in the prompt without classification. Table 1 shows that configuration (1) achieves the best performance and is even slightly better than using human-expert manually assigned classes, suggesting that the gpt-oss-based classifier is well aligned with the error patterns from our analysis. Configurations (2) and (4) perform similarly and both substantially outperform random classification in (3), indicating that our class structure and hints capture meaningful error modes, and that targeted hint selection remains helpful and robust even when classification is imperfect.
Table 1: Ablation on classification and hinting strategies on Mamo-Complex. We report single-turn and 5-turn accuracies (in percentage points).
Classification / Hint Strategy 1 turn 5 turn
gpt-oss classifier (ours) 83.81 89.90
Human expert manual classification 81.04 89.61
All hints, no classification 78.57 89.23
Random classes 72.57 87.23
No hints 70.00 86.57
6 Conclusion
We present OptiMind – a framework for formulating mixed-integer linear optimization problems with LLMs, combining a strong fine-tuned model with error-aware prompting and optional test-time scaling. Both training and inference benefit from domain-specific targeted hints that capture common error patterns within each optimization class and provide concise guidance to prevent those errors. The resulting pipeline attains strong performance across multiple benchmarks. While future LLMs may embed more expert knowledge, we believe the principles and techniques of our framework will remain essential, and we are keen to apply them to domains such as supply chain management or adapt them to enterprise-specific scenarios to drive real-world impact.
Appendix A Appendix
A.1 Additional Evaluation Results
Table 2: Accuracies of all models on expert-cleaned benchmarks in the single-sample setting (majority $K=1$) under different hint configurations.
@@ A.1 Additional Evaluation Results
Model Hints 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns
gpt-5 no $82.72\pm 2.15$ $83.23\pm 2.29$ $87.86\pm 1.55$ $89.36\pm 1.92$ $91.00\pm 1.32$ $92.79\pm 0.89$
yes $83.63\pm 2.17$ $85.15\pm 1.65$ $92.57\pm 1.19$ $93.42\pm 0.62$ $91.08\pm 1.76$ $92.79\pm 0.63$
o4-mini no $72.42\pm 2.33$ $78.58\pm 2.72$ $75.95\pm 2.38$ $87.78\pm 1.15$ $78.52\pm 2.66$ $88.37\pm 2.58$
yes $77.57\pm 2.55$ $0.30\pm 1.97 88$$8.09\pm 2.02$ $91.67\pm 1.52$ $84.18\pm 2.34$ $92.40\pm 1.49$
Qwen3-32B no $2.62\pm 2.73 66$$16\pm 1.74 66.$$6.67\pm 2.67$ $72.47\pm 2.65$ $75.78\pm 0.78$ $76.56\pm 3.12$
yes $68.18\pm 2.67$ $72.47\pm 2.65$ $76.78\pm 1.05$ $84.05\pm 1.48$ $76.30\pm 0.45$ $80.98\pm 0.45$
SIRL-Gurobi32B no $49.29\pm 3.39$ $51.81\pm 2.61$ $63.81\pm 1.53$ $74.44\pm 1.66$ $69.19\pm 1.48$ $74.10\pm 1.05$
yes $2.75\pm 2.18 55$$5.44\pm 2.04$ $4.33\pm 1.65 77$$7.72\pm 1.57$ $2.65\pm 1.00 77$$7.56\pm 1.67$
gpt-oss-20b no $66.23\pm 3.77$ $71.89\pm 3.05$ $65.03\pm 3.91$ $86.80\pm 1.33$ $74.87\pm 3.40$ $86.45\pm 2.13$
yes $68.28\pm 2.61$ $75.05\pm 1.84$ $75.85\pm 2,74$ $89.47\pm 1.27$ $76.95\pm 3.02$ $87.57\pm 1.24$
Our SFT no $68.97\pm 2.15$ $72.43\pm 3.43$ $85.77\pm 2.19$ $90.71\pm 1.24$ $\bf{84.82\pm 2.95}$ $\bf{89.62\pm 2.80}$
yes $\bf{70.50\pm 3.29}$ $\bf{75.25\pm 2.57}$ $\bf{89.09\pm 1.27}$ $\bf{92.04\pm 1.05}$ $2.34\pm 1.81 88$$8.35\pm 1.58$
Table 3: Accuracies under different base models and hint settings on benchmarks cleaned by parallel works. SIRL stands for Chen et al. (2025), LogiOR stands for Yang et al. (2025a), and Survey stands for Xiao et al. (2025a).
IndustryOR-SIRL IndustryOR-LogiOR IndustryOR-Survey MAMO-Complex-SIRL MAMO-Complex-Survey
Model K Hints 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns
gpt-5 1 no 69.67 71 68.29 70.12 66.67 68.25 72.57 74.05 65.76 66.21
yes 72 73 76.82 78.04 70.23 69.04 88.17 89.65 85.58 86.48
o4-mini 1 no 61.33 64.67 66.26 74.79 68.25 69.84 57.96 69.95 58.55 66.67
yes 68.67 70 71.54 74.39 66.67 63.49 81.77 86.53 78.82 86.48
SIRL-Gurobi32B 1 no 29.25 29.87 38.9 39 47.85 47.85 53.05 53.1 83.06 82.97
yes 30.8 31.2 35.1 35.6 46.67 46.67 65.56 65.61 77.02 77.2
gpt-oss-20b 1 no 54.3 62.7 59.87 65.97 62.85 67.38 60.39 77.83 68.1 78.82
yes 57.3 63.3 66.09 70 62.38 67.38 73.59 85.27 77.47 85.85
Our SFT 1 no 57.3 63.7 63.71 67.68 67.61 68.57 79.8 85.32 85.22 85.94
yes 59.2 62.6 67.68 71.64 69.04 71.42 84.43 87.68 87.38 90.09
Figure 11: Accuracy over five self-correcting turns on the three benchmarks.
A.2 Sample Problem Description and Output Optimization Model